OpenStudy (anonymous):

Let Z be a standard Normal distribution, w = z^2, find the moment generating function of W

OpenStudy (anonymous):

please help and explain

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how do i apply this i am new to this?

OpenStudy (anonymous):

We'll probably want the continuous integral rather than a discrete sum...

OpenStudy (perl):

$$ \Large{ M_x(t) = E(e^{tx})= \int e^{tx} f(x) } $$

OpenStudy (anonymous):

what is that \[M _{x} (t) = \int\limits_{-\infty}^{\infty} e ^{tx} f(x) dx\]

OpenStudy (anonymous):

ok

OpenStudy (perl):

now you can plug in \( \Large f(x) = \frac{1}{\sigma\sqrt{2\pi}}\, e^{-\frac{(x - \mu)^2}{2 \sigma^2}} \)

OpenStudy (perl):

lets use the variable z

OpenStudy (anonymous):

okay I'm following so far

OpenStudy (perl):

$$ \large { \text{Let Z be a random variable with mean = 0 , st. dev = 1 } \\ \Large M_z(t) = E(e^{tZ^2})= \frac{1}{\sqrt{2\pi}}\int_{-\infty }^{ \infty } e^{tz^2}e^{-\frac12 z^2} ~dz } $$

OpenStudy (anonymous):

okay so we plugged in for x with the z variable

OpenStudy (perl):

right

OpenStudy (perl):

and plug in mu = 0, sigma = 1 into f(x) above , to get the simplified expression

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can some one help me finish it after i plug in the mean and std. deviation is that it?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Recall the distribution function for the normally distributed random variable \(X\) with mean \(\mu\) and variance \(\sigma^2\): \[f_X(x)=\frac{1}{\sqrt{2\pi\sigma^2}}\exp\left(-\frac{(x-\mu)^2}{2\sigma^2}\right)\] where \(\exp(\cdots)=e^{\cdots}\), in case you're unfamiliar with the notation. For a standard normal distribution, we set \(\mu=0\) and \(\sigma^2=1\), call it \(Z\), and so \[f_Z(z)=\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{z^2}{2}\right)\] By the definition of the moment generating function, you have \[\begin{align*} \mathrm{M}_W(t)&=\mathbb{E}\left(\exp(tZ^2)\right)\\\\ &=\int_{-\infty}^\infty \exp(tz^2)\left(\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{z^2}{2}\right)\right)\,dz\\\\ &=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^\infty \exp\left(\left(t-\frac{1}{2}\right)z^2\right)\,dz\\\\ &=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^\infty \exp\left(-\frac{1}{2}\left(1-2t\right)z^2\right)\,dz\\\\ &=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^\infty \exp\left(-\frac{z^2}{2\left((1-2t)^{-1/2}\right)^2}\right)\,dz \end{align*}\] Nearly done. Let's substitution \(\xi=\dfrac{z}{(1-2t)^{-1/2}}\), then \(d\xi=\dfrac{dz}{(1-2t)^{-1/2}}\), and so the change of variables gives us the equivalent integral, \[\frac{\color{red}{(1-2t)^{-1/2}}}{\sqrt{2\pi}}\int_{-\infty}^\infty \exp\left(-\frac{\xi^2}{2}\right)\,d\xi\] What can you say about the integral above, not colored red?

OpenStudy (anonymous):

They cancel out ? @sithsandgiggles

OpenStudy (anonymous):

Let's be more careful with what you're trying to say... What exactly "cancels"? And what do you mean by "cancel"?

OpenStudy (anonymous):

they have similar terms? i guess thats what i noticed

OpenStudy (anonymous):

Here's a hint: we know that \[f_Z(z)=\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{z^2}{2}\right)\] is the standard normal distribution's density function defined over \(-\infty<z<\infty\). Replacing \(z\) with \(\xi\) doesn't change that fact. The definite integral over \((-\infty,\infty)\) is the total area under the curve (the density function above). What's the total area (i.e. probability)?

OpenStudy (anonymous):

is it 1?

OpenStudy (anonymous):

Right! This means \[\frac{\color{red}{(1-2t)^{-1/2}}}{\sqrt{2\pi}}\int_{-\infty}^\infty \exp\left(-\frac{\xi^2}{2}\right)\,d\xi=\color{red}{(1-2t)^{-1/2}}=\frac{1}{\sqrt{1-2t}}\] and this is the mgf for \(W=Z^2\).

OpenStudy (anonymous):

wow awesome ! great explanation

OpenStudy (zarkon):

you should also list the values of t for which this result holds.